# Introducing Unbiased Depth into 2D Gaussian Splatting for High-accuracy Surface Reconstruction

### v1

@@ Abstract
Recently, 2D Gaussian Splatting (2DGS) has demonstrated superior geometry reconstruction quality than the popular 3DGS by using 2D surfels to approximate thin surfaces. However, it falls short when dealing with glossy surfacess, resulting in visible holes in these areas. We found the reflection discontinuity causes the issue. To fit the jump from diffuse to specular reflection at different viewing angles, depth bias is introduced in the optimized Gaussian primitives. To address that, we first replace the depth distortion loss in 2DGS with a novel depth convergence loss, which imposes a strong constraint on depth continuity. Then, we rectified the depth criterion in determining the actual surface, which fully accounts for all the intersecting Gaussians along the ray. Qualitative and quantitative evaluations across various datasets reveal that our method significantly improves reconstruction quality, with more complete and accurate surfaces than 2DGS.
@@ 1 Introduction
During the past years, differentiable multi-view 3D reconstruction approaches [31, 10, 25, 27, 1, 26] have rapidly emerged along with the neural network revolution. Among these, 3D Gaussian Splatting (3DGS) [15] is recognized as one of the most appealing methods. It represents scenes with a collection of 3D Gaussian primitives, achieving remarkable performance in real-time photorealistic novel view synthesis. Since this innovation, follow-up works have flourished in a short period, advancing 3DGS in such as rendering [24, 22, 35], dynamic or large-scale scenes [28, 19, 12, 20, 16], and various other applications [8, 29, 30].
However, direct geometry extraction using the resulting Gaussian centers only leads to inaccurate reconstruction due to the volumetric property of 3D Gaussians, which conflicts with the thin surface nature. To address this issue, SuGaR [11] aligns Gaussians to surfaces via regularization, using depth maps to extract surface points. Other works [5, 4] attempt to flatten the 3D Gaussians. Recently, GOF [35] proposed an opacity field to derive the isosurface. Nevertheless, 3D Gaussians still retain thickness.
The recent work 2D Gaussian Splatting (2DGS) [13] makes a breakthrough, where 2D surfels are used instead of 3D Gaussians. 2D disks can align well with complex thin surfaces. Meanwhile, using 2D splats overcomes the inconsistency issue of 3D Gaussian primitives in multi-view projection, thereby significantly improving the geometric accuracy in reconstruction. Although theoretically sound, surface holes are often observed in 2DGS results, which arouses our curiosity; see, e.g., the 2nd column of Fig. LABEL:fig:teaser.
@@
In this paper, we propose an unbiased 2D Gaussian Splatting to address the discontinuity issue and depth bias mentioned above, with two key components. First, we propose a depth convergence loss, which forces the Gaussian depth to be continuous and smooth. Given the reliable depth around the highlight regions, the proposed constraint will promote an “edge-growing” effect from the specular-diffuse boundary to recover complete surface geometry for these areas. However, the highlight Gaussians will remain behind the surface using only the convergence loss. Therefore, we introduce depth correction, a new criterion to identify the actual surface, which considers both the number of intersecting Gaussians and the accumulated opacity along the ray. With unbiased depth, our improved 2DGS substantially enhances reconstruction quality. Experiments on various datasets have shown significant superiority of our method in geometric accuracy over the original 2DGS.
In summary, our contributions are as follows:
  • We analyze the root cause of holes when dealing with glossy surfaces in 2D Gaussian Splatting.
  • We propose a novel depth convergence loss that promotes “edge-growing” for high-specular regions, enabling complete and smooth surface reconstruction.
@@ 2 Related work
Recovering 3D geometry from 2D images is a fundamental yet challenging task in computer vision, including both traditional methods [1, 14, 33] and deep-learning-based methods [8, 15, 38]. Recent advances in neural radiance fields bring new insights to multi-view-based 3D reconstruction. This section mainly reviews recent work on neural implicit surface reconstruction and Gaussian splatting methods.
Neural Implicit Methods.
@@
Neural radiance fields (NeRF) [25] have achieved remarkable results in novel view synthesis (NVS) and have opened up new avenues for multi-view reconstruction. NeuS [27] and VolSDF [32] map volume density to signed distance fields (SDF), allowing mesh extraction at any resolution using the Marching Cubes [21]. NeuralWarp [7] enhances photometric consistency through a warping loss, improving reconstruction quality. Subsequent works [9, 34] leverage geometry priors, like normal priors and depth priors, have further refined reconstruction accuracy. Neuralangelo [18] employs hash encodings [26], a more expressive scene representation, achieving highly detailed reconstructions. Although these methods yield promising surface reconstruction, the long training time remains a significant bottleneck.
Gaussian Splatting.
@@
Recently, 3D Gaussian Splatting (3DGS) [15] has become one of the most popular 3D reconstruction methods, as it achieves high-quality real-time photorealistic NVS results. Follow-up works apply or improve 3DGS in various aspects, including rendering [24, 22, 35], reconstructing dynamic or large-scale scenes [28, 19, 12, 20, 16], and various other applications [8, 29, 30]. However, a remaining challenge is to derive high-accuracy geometry from the optimized Gaussian primitives.
@@
SuGaR [11] first addressed the geometry problem of 3DGS by regularizing the Gaussians to conform to surfaces. However, due to the inherent thickness of Gaussians, the extracted surface points are often noisy. Some recent works [5, 23, 33, 37] turn to SDF and jointly optimize 3DGS and neural SDF for surface reconstruction. GOF [35] constructs a Gaussian opacity field upon 3D Gaussians, identifying level sets to determine surfaces. Spiking GS [36] uses a minimal number of Gaussians, striking a good balance between geometric quality and cost. To approximate the thin nature of surfaces, NeuSG [5] proposes to minimize the minimal axis of 3D Gaussians to be flat shapes. PSGR [4] further incorporates multi-view geometric consistency, reaching remarkable reconstruction quality. Nonetheless, the planar Gaussians adopted in these methods are still 3D primitives with thickness.
@@
To acquire real 2D splats, Gaussian Surfels [6] directly zeroes the third axis of 3D Gaussians, with other components of 3DGS unchanged. More recently, 2DGS [13] rebuilt the Gaussian Splatting system using pure 2D disks, which completely solves the inconsistency of multi-view projections brought by 3D Gaussians, achieving very high reconstruction quality. However, 2DGS encounters depth bias in glossy areas, resulting in holes or limited geometric details. Our method addresses 2DGS’s limitations and introduces unbiased depth for surface reconstruction, significantly improving the geometry quality.
@@ 3.1 Preliminaries
3D Gaussian Splatting
represents 3D scenes with a set of 3D Gaussian primitives [15]. Each primitive is parameterized via a center point ${\mathbf{p}}_{k}$ and a covariance matrix $\boldsymbol{\Sigma}$:
@@
where $\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$, with $\mathbf{R}$ representing the rotation matrix and $\mathbf{S}$ as the scaling matrix. 3D Gaussians are transformed from the world coordinates to the camera coordinates using the transformation matrix $\mathbf{W}$. An affine transformation $\mathbf{J}$ [38] is applied to the projection to the image plane, resulting in the transformed covariance matrix:
@@
The first two rows and columns of $\boldsymbol{\Sigma}^{\prime}$ are extracted to derive a 2D Gaussian value $\mathcal{G}^{2D}(\mathbf{x})$, where $\mathbf{x}=(x,y)$ denotes an image coordinate. 3DGS [15] employs $\mathcal{G}^{2D}(\mathbf{x})$ for front-to-back volumetric alpha blending:
@@
2D Gaussian Splatting
uses 2D surfels to represent 3D scenes [13]. A 2D splat is composed of a center $\mathbf{p}_{k}$, two principal tangential vectors $\mathbf{t}_{u}$ and $\mathbf{t}_{v}$, and a scaling vector $\mathbf{S}=(s_{u},s_{v})$. By computing the ray-splat intersection $\mathbf{u}(\mathbf{x})$ in $uv$ space and applying an object-space low-pass filter [3], its 2D Gaussian value can be evaluated by:
@@
where $\mathbf{c}$ denotes the image coordinate of the projection of $\mathbf{p}_{k}$, and $\sigma$ is the radius of screen-space low-pass filter. $\sigma$ is set to 2/$\sqrt{2}/2$ to handle degenerate cases when 2D Gaussian is rasterized from a slanted viewing angle.
2DGS uses a rasterization process similar to that of 3DGS [15]. Initially, screen-space bounding boxes are computed for each Gaussian primitive. Then, 2D Gaussians are sorted by depth and organized into tiles based on their bounding boxes. Finally, volumetric alpha blending is performed from front to back until the accumulated opacity reaches saturation, which can be formulated as:
@@ 3.2 Hole Formation in Glossy Areas
Fig. 2 visualizes a real case of the resulting Gaussians from 2DGS. It clearly shows that in glossy areas, only a few transparent (low-opacity) Gaussian primitives are left at the actual surface, whereas most splats with much greater opacity concave inward after the optimization. We further attribute the issue to the depth distortion loss proposed in 2DGS [13], where the intersected primitives along a ray are constrained to be close to each other using this loss. However, the multiplication of opacity as weights makes those splats with higher opacity to dominate the loss calculation (see Eq. 13 in their paper). Therefore, in regions of specular highlights, it will lead to the clustering of Gaussian surfels towards the inward-concave, forming a “false” surface. When extracting the object mesh, the algorithm fetches biased depth in these areas, resulting in holes or pits.
@@ 3.3 Depth Convergence Loss
As discussed, in the glossy regions of the object, the surface contains a sparse distribution of low-opacity Gaussians. To improve Gaussian coveragence on the actual surface without adversely impacting other diffuse surfaces, we encourage Gaussians to move toward the actual surface during optimization, thereby enhancing reconstruction quality. Accordingly, we introduce a new depth convergence loss, and use it to replace the depth distortion loss in 2DGS [13].
@@
To further encourage concave Gaussians to converge toward the actual surface, we scale the partial derivative of $D_{i}$ w.r.t. $d_{i}$ by a factor greater than 1 during gradient calculation: $\frac{\partial}{\partial d_{i}}D_{i}\leftarrow 2k(d_{i}-d_{i-1})$, where $k=1.25$.
@@
Fig. 3 visualizes the process of hole filling brought by our depth convergence loss. In regions with holes, the loss gradually pulls concave Gaussians from the hole margins to the true surface, exhibiting an “edge-growing” phenomenon along with optimization. Our loss also induces larger gradients for Gaussians in glossy areas, which might further trigger the Gaussian splitting and cloning mechanism [15], thereby generating additional Gaussians to fill the holes.
@@ 3.4 Depth Correction
When $T_{i}$ decreases to 0.5, nearby depth points are employed to compute the surface normal with finite differences [13].
@@
where $\epsilon$ is a small constant compensating opacity, and can be used to control the impact of the cumulative number of intersected Gaussians on $O_{i}$. We set $\epsilon$ as 0.1 in all our experiments. When the accumulated opacity exceeds a threshold (0.6 according to experiments), the intersection point $\mathbf{p}_{s}$ is considered as the actual surface point of the object. See our supple material for more analysis of depth correction.
Next, we derive the normal map from the gradients of the rectified depth map and align it with the splats’ normal [13]:
@@
where $\mathbf{N}$ represents the normal estimated at the surface point, $\mathbf{n}_{i}$ represents the normal of the $i$-th splat that is oriented towards the camera, and $w_{i}$ denotes the blending weight of the $i$-th intersetion point.
@@ 3.5 Final Loss
Following 3DGS [15], we combine $\mathcal{L}_{1}$ loss with D-SSIM as the color reconstruction loss. Inspired by PGSR [4], we applied gamma correction with a gamma value of $0.5$ to the ground truth images to enhance contrast in dark areas, and then calculated the color loss with the rendered image:

### v3

@@ Abstract
Recently, 2D Gaussian Splatting (2DGS) has demonstrated superior geometry reconstruction quality than the popular 3DGS by using 2D surfels to approximate thin surfaces. However, it falls short when dealing with glossy surfaces, resulting in visible holes in these areas. We find that the reflection discontinuity causes the issue. To fit the jump from diffuse to specular reflection at different viewing angles, depth bias is introduced in the optimized Gaussian primitives. To address that, we first replace the depth distortion loss in 2DGS with a novel depth convergence loss, which imposes a strong constraint on depth continuity. Then, we rectify the depth criterion in determining the actual surface, which fully accounts for all the intersecting Gaussians along the ray. Qualitative and quantitative evaluations across various datasets reveal that our method significantly improves reconstruction quality, with more complete and accurate surfaces than 2DGS. Code is available at https://github.com/XiaoXinyyx/Unbiased_Surfel.
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010372</concept_id> <concept_desc>Computing methodologies Rendering</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293</concept_id> <concept_desc>Computing methodologies Machine learning approaches</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010396</concept_id> <concept_desc>Computing methodologies Shape modeling</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
[…]
@@ 1 Introduction
During the past years, differentiable multi-view 3D reconstruction approaches [YLL18, GFZ19, MST21, WLL21, BMT21, MESK22] have rapidly emerged along with the neural network revolution. Among these, 3D Gaussian Splatting (3DGS) [KKLD23] is recognized as one of the most appealing methods. It represents scenes with a collection of 3D Gaussian primitives, achieving remarkable performance in real-time photorealistic novel view synthesis. Since this innovation, follow-up works have flourished in a short period, advancing 3DGS in such as rendering [MST23, LYX24, YSG24], dynamic or large-scale scenes [WYF24, LKL23, GZL24, LLT24, KMK24], and various other applications [DCZ24, XHL24, YLF24].
However, direct geometry extraction using the resulting Gaussian centers only leads to inaccurate reconstruction due to the volumetric property of 3D Gaussians, which conflicts with the thin surface nature. To address this issue, SuGaR [GL23] aligns Gaussians to surfaces via regularization, using depth maps to extract surface points. Other works [CLL23, CLY24] attempt to flatten the 3D Gaussians during optimization. Recently, GOF [YSG24] proposed an opacity field to derive the isosurface. Nevertheless, 3D Gaussians still retain thickness.
The recent work 2D Gaussian Splatting (2DGS) [HYC24] makes a breakthrough, where 2D surfels are used instead of 3D Gaussians. 2D disks can align well with complex thin surfaces. Meanwhile, using 2D splats overcomes the inconsistency issue of 3D Gaussian primitives in multi-view projection, thereby significantly improving the geometric accuracy in reconstruction. Although theoretically sound, surface holes are often observed in 2DGS results, which arouses our curiosity; see, e.g., the 2nd column of Fig. Introducing Unbiased Depth into 2D Gaussian Splatting for High-accuracy Surface Reconstruction.
@@
In this paper, we propose an unbiased 2D Gaussian Splatting to address the discontinuity issue and depth bias mentioned above, with two key components. First, we propose a depth convergence loss, which forces the Gaussian depth to be continuous and smooth. Given the reliable depth around the highlight regions, the proposed constraint will promote an “edge-growing” effect from the specular-diffuse boundary to recover complete surface geometry for these areas. However, the highlight Gaussians will remain behind the surface using only the convergence loss. Therefore, we introduce depth correction, a new criterion to identify the actual surface, which considers both the number of intersecting Gaussians and the accumulated opacity along the ray. With unbiased depth, our improved 2DGS substantially enhances reconstruction quality. Experiments on various datasets show that our method significantly outperforms the original 2DGS in terms of geometric accuracy.
In summary, based on experimental analysis of glossy regions, we present the following contributions:
  • We propose a novel depth convergence loss to enhance stronger geometric constraints, thereby promoting “edge-growing” for high-specular regions, leading to more complete and smoother surface reconstruction than the original 2DGS.
@@ 2 Related work
Recovering 3D geometry from 2D images is a fundamental yet challenging task, including both traditional methods [1, 14, 33] and deep-learning-based methods [8, 15, 38]. Recent advances in neural radiance fields bring new insights to multi-view-based 3D reconstruction. This section mainly reviews recent work on neural implicit surface reconstruction and Gaussian splatting methods.
@@
Neural Implicit Methods. Neural radiance fields (NeRF) [MST21] have achieved remarkable results in novel view synthesis (NVS) and have opened up new avenues for multi-view reconstruction. NeuS [WLL21] and VolSDF [YGKL21] map volume density to signed distance fields (SDF), allowing mesh extraction at any resolution using the Marching Cubes [LC97]. NeuralWarp [DBD22] enhances photometric consistency through a warping loss, improving reconstruction quality. Subsequent works [FXOT22, YPN22] leverage geometry priors, like normal priors and depth priors, have further refined reconstruction accuracy. Neuralangelo [LME23] employs hash encodings [MESK22], a more expressive scene representation, achieving highly detailed reconstructions. Although these methods yield promising surface reconstruction, the long training time remains a significant bottleneck.
@@
Gaussian Splatting. Recently, 3D Gaussian Splatting (3DGS) [KKLD23] has become one of the most popular 3D reconstruction methods, as it achieves high-quality real-time photorealistic NVS results. Follow-up works apply or improve 3DGS in various aspects, including rendering [MST23, LYX24, YSG24], exploring new shading methodologies for specular surfaces  [ZCW25, JTL23, KQK24, TC24], reconstructing dynamic or large-scale scenes [WYF24, LKL23, GZL24, LLT24, KMK24], and various other applications [DCZ24, XHL24, YLF24]. However, a remaining challenge is to derive high-accuracy geometry from the optimized Gaussian primitives.
@@
SuGaR [GL23] first addressed the geometry problem of 3DGS by regularizing the Gaussians to conform to surfaces. However, due to the inherent thickness of Gaussians, the extracted surface points are often noisy. Some recent works [CLL23, LSH24, YLX24, ZLH24] turn to SDF and jointly optimize 3DGS and neural SDF for surface reconstruction. GOF [YSG24] constructs a Gaussian opacity field upon 3D Gaussians, identifying level sets to determine surfaces. Spiking GS [ZLM24] uses a minimal number of Gaussians, striking a good balance between geometric quality and cost. To approximate the thin nature of surfaces, NeuSG [CLL23] proposes to minimize the minimal axis of 3D Gaussians to be flat shapes. PSGR [CLY24] further incorporates multi-view geometric consistency, reaching remarkable reconstruction quality. Nonetheless, the planar Gaussians adopted in these methods are still 3D primitives with thickness.
@@
To acquire real 2D splats, Gaussian Surfels [DXX24] directly zeroes the third axis of 3D Gaussians, with other components of 3DGS unchanged. 2DGS [HYC24] rebuilt the Gaussian Splatting system using pure 2D disks, which completely solves the inconsistency of multi-view projections brought by 3D Gaussians, achieving very high reconstruction quality. However, 2DGS encounters depth bias in glossy areas, resulting in holes or limited geometric details. Our method addresses 2DGS’s limitations and introduces unbiased depth for surface reconstruction, significantly improving the geometry quality.
@@ 3.1 Preliminaries
3D Gaussian Splatting represents 3D scenes with a set of 3D Gaussian primitives [KKLD23]. Each primitive is parameterized via a center point ${\mathbf{p}}_{k}$ and a covariance matrix $\boldsymbol{\Sigma}$:
@@
where $\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$, with $\mathbf{R}$ representing the rotation matrix and $\mathbf{S}$ as the scaling matrix. 3D Gaussians are transformed from the world coordinates to the camera coordinates using the transformation matrix $\mathbf{W}$. An affine transformation $\mathbf{J}$ [ZPVBG01] is applied to the projection to the image plane, resulting in the transformed covariance matrix:
@@
The first two rows and columns of $\boldsymbol{\Sigma}^{\prime}$ are extracted to derive a 2D Gaussian value $\mathcal{G}^{2D}(\mathbf{x})$, where $\mathbf{x}=(x,y)$ denotes an image coordinate. 3DGS [KKLD23] employs $\mathcal{G}^{2D}(\mathbf{x})$ for front-to-back volumetric alpha blending:
@@
2D Gaussian Splatting uses 2D surfels to represent 3D scenes [HYC24]. A 2D splat is composed of a center $\mathbf{p}_{k}$, two principal tangential vectors $\mathbf{t}_{u}$ and $\mathbf{t}_{v}$, and a scaling vector $\mathbf{S}=(s_{u},s_{v})$. By computing the ray-splat intersection $\mathbf{u}(\mathbf{x})$ in $uv$ space and applying an object-space low-pass filter [BHZK05], its 2D Gaussian value can be evaluated by:
@@
where $\mathbf{c}$ denotes the image coordinate of the projection of $\mathbf{p}_{k}$, and $\sigma$ is the radius of screen-space low-pass filter. $\sigma$ is set to $\sqrt{2}/2$ to handle degenerate cases when 2D Gaussian is rasterized from a slanted viewing angle.
2DGS uses a rasterization process similar to that of 3DGS [KKLD23]. Initially, screen-space bounding boxes are computed for each Gaussian primitive. Then, 2D Gaussians are sorted by depth and organized into tiles based on their bounding boxes. Finally, volumetric alpha blending is performed from front to back until the accumulated opacity reaches saturation, which can be formulated as:
@@ 3.2 Hole Formation in Glossy Areas
Fig. 2 visualizes a real case of the resulting Gaussians from 2DGS. It clearly shows that in glossy areas, only a few transparent (low-opacity) Gaussian primitives are left at the actual surface, whereas most splats with much greater opacity concave inward after the optimization. We further attribute the issue to the depth distortion loss proposed in 2DGS [HYC24], where the intersected primitives along a ray are constrained to be close to each other using this loss. However, the multiplication of opacity as weights makes those splats with higher opacity dominate the loss calculation (see Eq. 13 in their paper). Therefore, in regions of specular highlights, it will lead to the clustering of Gaussian surfels towards the inward-concave, forming a “false” surface. When extracting the object mesh, the algorithm fetches biased depth in these areas, resulting in holes or pits.
@@ 3.3 Depth Convergence Loss
As discussed, in the glossy regions of the object, the surface contains a sparse distribution of low-opacity Gaussians. To improve Gaussian coveragence on the actual surface without adversely impacting other diffuse surfaces, we encourage Gaussians to move toward the actual surface during optimization, thereby enhancing reconstruction quality. Accordingly, we introduce a new depth convergence loss, and use it to replace the depth distortion loss in 2DGS [HYC24].
@@
To further encourage concave Gaussians to converge toward the actual surface, we scale the partial derivative of $D_{i}$ w.r.t. $d_{i}$ by a factor greater than 1 during gradient calculation: $\frac{\partial}{\partial d_{i}}D_{i}\leftarrow 2k(d_{i}-d_{i-1})$, where $k=1.25$. This modification encourages Gaussians to shift toward the camera direction. For concave highlight regions where multi-view consistency constraints prove less effective, our loss function progressively reduces concavity depth. In strongly textured regions, multi-view consistency constraints dominate the optimization, while the cconvergence loss primarily serves to aggregate Gaussians.
@@
Fig. 3 visualizes the process of hole filling brought by our depth convergence loss. In regions with holes, the loss gradually pulls concave Gaussians from the hole margins to the true surface, exhibiting an “edge-growing” phenomenon along with optimization. Our loss also induces larger gradients for Gaussians in glossy areas, which might further trigger the Gaussian splitting and cloning mechanism [KKLD23], thereby generating additional Gaussians to fill the holes.
@@ 3.4 Depth Correction
When $T_{i}$ decreases to 0.5, nearby depth points are employed to compute the surface normal with finite differences [HYC24].
@@
where $\epsilon$ is a small constant compensating opacity, and can be used to control the impact of the cumulative number of intersected Gaussians on $O_{i}$. We set $\epsilon$ as 0.1 in all our experiments. When the accumulated opacity exceeds a threshold (0.6 according to experiments), the intersection point $\mathbf{p}_{s}$ is considered as the actual surface point of the object. See our supplementary material for more analysis of depth correction.
Next, we derive the normal map from the gradients of the rectified depth map and align it with the splats’ normal [HYC24]:
@@
where $\mathbf{N}$ is the normal estimated at the surface point, $\mathbf{n}_{i}$ represents the normal of the $i$-th splat that is oriented towards the camera, and $w_{i}$ denotes the blending weight of the $i$-th intersetion point.
@@ 3.5 Final Loss
Following 3DGS [KKLD23], we combine $\mathcal{L}_{1}$ loss with D-SSIM as the color reconstruction loss. Inspired by PGSR [CLY24], we applied gamma correction with a gamma value of $0.5$ to the ground truth images to enhance contrast in dark areas, and then calculated the color loss with the rendered image: